Fifteen children’s books cost $51.25. Some were priced at $2.25 each, and the rest of the books were sold at $4.75 each. How many books were purchased at each price?
Answer by ikleyn(52781)

.
Fifteen children’s books cost $51.25. Some were priced at $2.25 each,
and the rest of the books were sold at $4.75 each.
How many books were purchased at each price?
~~~~~~~~~~~~~~~

x books at the price $4.75 each and

(15-x) books at the price of $2.25 each.


Write the total money equation

    4.75x + 2.25*(15-x) = 51.25  dollars.


Simplify and find x

    4.75x + 2.25*15 - 2.25x = 51.25

    4.75x - 2.25x = 51.25 - 2.25*15

       2.50x      =     17.50

           x      =     17.50/2.50 = 7.


ANSWER.  7 books at $4.75 each, and the rest 15-7 = 8 books at $2.25 each.


CHECK total money:  7*4.75 + 8*2.25 = 51.25  dollars.    ! Correct !

Solved.